# Explainable Artificial Intelligence for Autonomous Driving: A Comprehensive Overview and Field Guide for Future Research Directions

### v4

@@ 3.1 The need for explanations in AVs
The need for explanations in autonomous driving arises from existing issues, established regulations and standards covered in previous subsections, and cross-disciplinary views and opinions of society. At the highest level, the necessity of explanations for AVs can be summarized in terms of four perspectives:
@@
Overall, we can conclude that the explainability of autonomous driving systems is an expectation and a requirement from a multidisciplinary point of view (Figure 3).
3.2 Potential benefits of explanations in AVs
@@
  • Human-Centered Design
    Getting the end users’ inputs, opinions, and anticipations on the design and development of the semi or fully AVs can help with the acceptance of this technology by the general community [73]. In this context, it is essential to provide user-friendly human-machine interfaces (HMIs) to the stakeholders, such as backup drivers and passengers. Studies show that explanations can be conveyed to end users via light, visual, audio, and textual information using supplied HMIs [74, 75, 76]. For instance, Schneider et al.’s recent empirical study confirms that applying stakeholders’ multi-modal feedback to the simulated design of autonomous driving creates a positive user experience [77]. Thus, a human-centered design meeting users’ needs and expectations is a necessary step for the widespread adoption of AVs technology.
  • Trustworthiness
    As careless and hazardous driving can directly impact the safety of passengers and bystanders, people naturally require an acceptable level of safety of AVs. In addition, understanding the causes of actions or decisions is a natural human requirement. As stated in Riberio et al.’s [78] work, “if the users do not trust a model or a prediction, they will not use it.” Holliday et al. [79] have also empirically shown that providing explanations and perceptible systems significantly increases users’ trust in a system. On the other hand, frequently occurring failures without faithful explanations to stakeholders can seriously damage individual and public trust in intelligent systems. Once trust in an intelligent system is damaged, regaining it can be onerous [80]. In this sense, Israelson and Ahmed [81] have shown that there is an inherent need for algorithmic assurance to build trust in human-autonomous system relationships in their analysis.
  • Traceability
    Explainable intelligent driving systems can help forensic analysts and system auditors understand the entire decision-making process of an autonomous car during the journey via a post-trip analysis. A recorded log data of action-explanation pairs can help understand the car’s temporal actions and shed light on whether particular temporal actions of the car were correct or incorrect. Hence, traceability is an essential requirement of self-driving and can help debug and analyze the existing driving system and provide an opportunity for further enhancements.
  • Transparency and accountability
    If trustworthiness is achieved in intelligent decision-making, it further brings transparency and accountability to AVs technology. Martinho et al. [38] have noted that accountability combines liability and responsibility as a broader concept. In the context of autonomous driving, accountability can be defined as compliance with established legal principles in specific jurisdictions. As the regulatory standards urge the “right to an explanation” as required by GDPR [49], accountability becomes a crucial concept that combines social expectations and legislative norms on the autonomous driving spectrum. In addition, achieving accountability also helps to deal with potential liability and responsibility gaps, as defined by Burton et al., [39], in potential post-accident investigations with the involvement of autonomous cars. Recently Mercedes-Benz has taken a promising step forward and announced that the corporation will take legal responsibility for any accidents that their self-driving systems are engaged in [82]. Mercedes’s declaration of legal culpability is a significant milestone toward the accountability of AVs technology. Hence, explanations promise immense benefits to road users, engaged stakeholders, and AVs manufacturers in general.
@@ 3.3 Explanation recipients in AVs
The details, types, and delivery of explanations vary in accordance with users’ identities, technical background knowledge in autonomous driving, and their various functional and cognitive abilities [67, 83]. For instance, a user having little technical expertise on how autonomous cars operate may be satisfied with a simple explanation of a relevant decision/outcome. However, an autonomous systems engineer will need more informative explanations to understand the current functionalities of the car, with the motivation to appropriately “debug” the existing driving system as required. Therefore, the use of domain knowledge and expertise of the explainee is essential to provide pertinent, sufficiently informative, and intelligible explanations [84, 85]. Motivated by a target audience definition of [86] and [67], we can distinguish four groups of the stakeholders in autonomous driving, namely Group 1 - Road users, Group 2 - AVs developers, Group 3 - Regulators and insurers, and Group 4 - Executive management of automobile companies. Figure 4 provides the identity of such stakeholders and their positions in the corresponding classification.
@@ 3.4 How to deliver explanations in AVs?
As explainees are classified based on their domain knowledge and needs, explanations and their design and evaluation techniques also vary depending on the context and knowledge of the category of explainees. In fact, explanation construction is one of the major challenges in current XAI research. Zablocki et al. [87] define four “W” questions in XAI-based autonomous driving: 1) Who needs explanations? 2) Why are explanations needed? 3) What kind of explanations can be generated? and 4) When should explanations be delivered? In general, explanations in AI can be distinguished based on their derivation category and classification. Some of the early practical studies have applied explanations to automated collaborative filtering systems [88] and knowledge-intensive case-based reasoning [89]. Another empirical approach attempted to derive explanations based on some intelligibility types [90] and used “why,” “why not,” “what if,” and “how to” type explanations for causality filtering. Furthermore, Liao et al. [91] have interviewed twenty user-interface and design practitioners working in different areas of AI to understand users’ explanatory requirements. By doing so, they have attempted to find the gaps in the interviewers’ products and developed a question bank: the authors represent users’ needs as questions so that users may potentially ask about the outcomes produced by an AI system. Overall, the stakeholder needs-based explanation design can be viewed as one of the promising approaches.
Another popular approach to producing explanations is based on using psychological tools from formal theories, according to the literature review of [92]. Depending on the context and addressee, both explanation derivation methods confirm their usefulness. These explanation generation approaches can find alignment in their application in autonomous driving; since autonomous driving involves people with
@@
diverse backgrounds in society, relevant XAI design needs inherent adjustments to the context problem. Like their derivation type, explanations also differ depending on the class in which they are included. Through an extensive survey, Omeiza et al. [67] propose the following dimensions of explanations in the context of autonomous driving:
Explanations based on cause filters: Based on available knowledge, explanations use predefined causes to explain the outcome of an event. The explanations are generated based on cause filters such as “why,” “why not,” “how to,” and “what if” queries (e.g., “Why did the car take the left lane instead of the right lane?”). In fact, we can note that this kind of explanations can be used across many autonomous driving operations.
Explanations based on content type: In this category, explanations are classified based on the components or elements involved with the explanations and the way they are presented. Examples of content types include input influence, input sensitivity, case bases, and demographic factors (for instance, explanations based on what input variables (i.e., driving features) contribute more to the predictive actions).
Explanations depending on a model: Here, explanations are distinguished by being either model agnostic or model dependant. For instance, some autonomous driving operations can be condition-specific and some can be general, regardless of the driving conditions (for example, explaining any autonomous driving action, as a model-agnostic rationale, can be specified for this group of explanations).
Explanations based on a system type: This category attempts to capture the properties of the operational system: Omeiza et al. [67] distinguish two kinds of explanations as either data-driven (i.e., explaining the outcome of a predictive model) or goal-driven (explaining an agent’s behaviors based on achieving its goal in a predefined setting).
Explanations with interactivity: Once an explanation is provided, a user may further ask follow-up questions to further understand a provided explanation. This feature brings interactivity into the explanation framework (e.g., a user interface for passengers on board that provides real-time explanations for corresponding actions).
Explanations with concrete scope: This category captures the feasibility and range of explanations that the system can generate by being either local or global. Local explanations are limited to explanations on some or a subset of all possible actions (i.e., explaining a single prediction in a specific traffic scenario). Global explanations, on the other hand, are capable of explaining all high-level decisions from an initial point to the destination, such as why an autonomous car chose a specific map, why it changed the
Figure 5: An example of a CNN for object classification in a real-time traffic scenario.
preplanned route during the travel, and so on.
@@
While Omeiza et al.’s [67] analysis of explanations above primarily focuses on informational content, the effective communication of explanations is also a key factor for good human-machine teaming. In general, the conveyance of explanations to end users is realized through a user interface (UX) or HMI of various types [93]. For instance, an HMI may be an interface to alert the human driver to take over the control of an autonomous vehicle in an emergent situation. Other potential examples are display monitors, sound, light signals, and vibrotactile technology that explain the vehicle’s decision-making intentions and bring situation awareness for people in the loop, as shown in Schneider et al.’s work [77].
4 AI for autonomous driving
Real-time decisions in autonomous driving are based on environmental perception, processing temporal sequence data, and mapping of real-time perception to relevant actions. In this regard, before reviewing relevant XAI techniques, we provide a brief overview of frontier AI approaches - convolutional neural networks and recurrent neural networks with their variations, reinforcement learning, and imitation learning. These AI techniques and their enhancements are dominant techniques that have proven empirical successes in the development of state-of-the-art AVs.
4.1 Convolutional neural networks
Convolutional neural networks (CNN) are an AI architecture to process spatial information, such as images and videos [94]. As a powerful learning technique, CNNs can detect discriminant visual features automatically from an input image and are extensively used for pattern recognition, object classification and detection, and other computer vision applications. CNNs can be regarded as universal nonlinear function approximators. The input x of each layer in a CNN model is organized in three dimensions: width, height, and depth. A typical convolutional neural network is parameterized by a weight vector consisting of a set of weights, W, between neurons and a set of bias values, b:
$\theta=[W,b]$ (1)
During training, a variety of useful and high-level features are extracted in the convolution layer. Then a pooling layer is used to reduce the size of the acquired feature map, to decrease computational costs. After that, the output of these steps is passed to the fully connected layer where neurons, along with the weights and biases, are connected with one another, and a nonlinear activation function is applied to the output of the previous step. Commonly used activation functions with CNN are Sigmoid, Tanh, and ReLu, while ReLu is the most preferred because of its relatively faster convergence. The network further makes
Figure 6: The network structure RNN and LSTM
a final prediction. In the context of autonomous driving, the role of CNN is indispensable for real-time scene understanding tasks, such as object detection, identification, segmentation, and classification. A typical example of CNN architecture for autonomous driving is shown in Figure  5.
While traditional CNNs have been successfully applied to numerous computer vision problems, a limitation of this learning architecture is that it can solely process Euclidean spatial data [95, 96]. However, in real-world applications, there might be a need to capture the spatial relationship between different elements necessitating the processing of non-Euclidean data. In this context, Graph CNNs (GCNNs) [97] have emerged as an alternative to conventional CNNs and are capable of processing any kind of non-grid data by representing elements in a graph structure. In autonomous driving, traffic data such as road networks, traffic flow, and spatial relationships among static and dynamic objects of such a network can be naturally represented as graphs, and GCNNs can effectively capture the dependencies and interactions among these elements. Moreover, as the driving environment is generally dynamic and rapidly changing, an inherent structure of GCNN can allow it to adapt to changes in the environment by dynamically updating the graph structure based on the provided sensor inputs and contextual information. Overall, both traditional CNNs and GCNNs and their augmented variations hold promise for dealing with a variety of challenging vision problems in autonomous driving.
4.2 Recurrent neural networks
Recurrent neural networks (RNNs) are a deep learning architecture for processing temporal and sequential data, such as time series, video, and natural language data [98, 99]. RNNs have a feedback loop to iterate over time phases of sequential data: the output of a previous time step becomes an input to the current step. While iterating over the different time steps, recurrent networks can maintain internal states that contain information about each time step; thus, RNN architectures leverage the concept of “memory,” which uses information from a previous input to yield output in the next time step. A typical RNN architecture has three layers: input, hidden, and output layers. The input layer consists of N units. A sequence of vectors of each time step t denoted as $\{...,{x}_{t-1},{x}_{t},{x}_{t+1},...\}$ is the input of this layer. The input layer is connected to a hidden layer where connections between the units are defined by means of a weight matrix. The hidden units of a hidden layer connect with each other through recurrent connections, and by such a structure, the hidden layer defines the memory of the entire network, formulated as
${h}_{t}=f_{H}({o}_{t}),$ (2)
in which
${o}_{t}={W}_{IH}{x}_{t}+{W}_{HH}{h}_{t-1}+{b}_{h}.$ (3)
The hidden layer is also connected to the output layer with weights ${W}_{HO}$, and based on such a network flow, the units of the output layer are calculated as follows:
${y}_{t}=f_{O}({W}_{HO}{h}_{t}+{b}_{o}).$ (4)
Figure 7: A diagram of reinforcement learning for a self-driving vehicle.
Similar to CNNs, $f_{O}(\cdot)$ is the hidden layer activation function, and ${b}_{h}$ is the bias vector of units of the hidden layer.
A major issue with traditional RNNs is the so-called vanishing gradient problem arising during training: network weights might not be effectively updated if the network is very deep. This may result in considerably small weight values that may reduce the network’s learning ability and force RNNs to have a limited memory capacity. To solve this problem, long short-term memory (LSTM) has been proposed as an enhancement of RNN, which can handle sequential data more effectively and learn better [100]. Compared to a simple RNN architecture, LSTM has “gates” that control the flow of information through the network. Having the ability to preserve long-term memory and retain earlier information in sequence makes LSTMs more practical than traditional RNNs. In general, both RNN and LSTM can be used to predict future position, velocity, and other parameters of autonomous driving. Figure 6 shows a structure of a typical RNN and LSTM.
4.3 Reinforcement learning
Reinforcement learning (RL) is a learning approach where an autonomous software agent interacts with an operational environment and learns to improve its performance via such interactions [101]. RL is a powerful machine learning framework for making real-time and sequential decisions. Generally, sequential decision-making problems are formalized within a setting formally known as Markov decision processes (MDP). An MDP comprises the following parameters:
  • S - a set of states
  • A - a set of actions
  • T- a transition function
  • R - a reward function
  • $\gamma$ a discount factor defined as a fixed value in the (0, 1] interval.
In such a setting, selecting an action ${a\in A}$ results in a new state $s\in S$ with a transition probability ${T(s,a,s^{\prime})\in(0,1)}$, and gives a reward ${R(s,a)}$ to an agent. The goal of an RL agent (i.e., a self-driving vehicle, in our case) is to discover the optimal policy $\pi^{*}$ that results in the maximum expected sum of discounted reward:
[…]
An agent’s reward in starting from a state s taking action a by following a policy $\pi$ is formulated as an action-value function (or Q-function) and defined as follows:
$Q_{\pi}(s,a)=\mathbb{E}_{\pi}\Bigg{\{}\sum_{k=0}^{H-1}\gamma^{k}r_{k+1}\mid s_{0}=s,a_{0}=a\Bigg{\}}.$ (6)
Basically, the value function is a measure of how good it is for an agent to be in a particular state [101]. The horizon H is the number of time steps in a given MDP. The MDP setting for RL uses the Markov property: the current state transition depends only on the previous state and previous action. But often, in real-world problems, such as in autonomous driving, an agent might not be able to capture all information and fail to observe all the states of the operational environment. In such a situation, an agent’s interaction with the surroundings is constructed as a partially observable MDP (POMDP). In the POMDP setting, states are replaced by observations [102]. Observations are generated by a latent state, which is not available to an agent in a POMDP. The natural state for a POMDP setting is the distribution on the latent state; this is called a belief state [101]. We can infer that a typical observation is considerably less informative than a natural Markov state, $S_{t}$. So, depending on how perception defines the traffic scenario and tasks, both MDP and POMDP can be used for sequential decision-making in autonomous driving. A general diagram of RL for autonomous driving is described in Figure 7.
4.4 Imitation learning
Imitation learning (IL) is a learning approach mimicking human behavior [103]. Also known as learning from demonstration, the agent learns from a dataset of demonstrations of human experts’ behavior in contrast to an RL agent. The ultimate goal of IL is that the agent imitates the actions of a human expert, and its learned policy $\pi_{il}$ matches the human expert’s policy $\pi_{h}$. There are two well-known types of IL:
1) Behaviour cloning (BC): In this category of IL, a model is trained in a supervised learning manner using a trajectory of state-action pairs $\{(s_{0},a_{0},s_{1},a_{1},\cdots)\}$. The goal is to minimize the loss function that calculates how the IL agent’s actions are different from an expert’s actions. The advantage of behaviour cloning is that is usually simple and efficient and does not require specific reward shaping. On the other hand, misleading correlations between input and output can affect the imitator’s learning ability seriously and lead to compounding errors.
2) Inverse Reinforcement Learning (IRL): Also referred to as Inverse Optimal Control (IOC), IRL learns an underlying reward function that the human expert aims to maximize. Once the agent infers the expert’s reward function, the optimal policy can then be learned via RL. Overall, empirical studies show that IRL can generalize better than BC in unseen states [104]; however, it is computationally costly and more intricate.
5 XAI for autonomous driving: A survey on state of the art
@@ 5.1 Previous surveys on XAI for AVs
There exist reviews on XAI for AVs, which provide valuable insights from various perspectives. These studies expose a variety of approaches, from a universal look to an algorithmic point of view. In this sense, the first noteworthy review on XAI for AVs is Omeiza et al.’s work [67]. They study the need for/role of explanations for autonomous driving, and focus on legal requirements, standards, and consumer expectations for the design and development of explainable autonomous driving systems. This provides their basis to present a conceptual XAI framework for modular autonomous driving. In further work, Zablocki et al. [87] present a detailed overview of end-to-end vision-based autonomous driving systems and describe explainability hurdles for AVs from an ML perspective. Finally, in very recent work, Kuznietsov et al. [105] present a systematic review of XAI techniques for modular and end-to-end autonomous driving by focusing on how such techniques improve safety and user trust. In this regard, they propose the SafeX framework for modular autonomous driving by integrating user interface, safety, and explainability.
Table 1: Reviews on XAI for AVs with their focus and coverage
Study Pipelines explored Proposed framework Form and content of explanations? Time granularity of explanations? Emerging paradigms? Methodological taxonomy? Omeiza et al., 2021 [67] modular modular Zablocki et al., 2022 [87] end-to-end Kuznietsov et al., 2024 [105] modular and end-to-end modular Our study modular and end-to-end end-to-end
Our study, as complement, extends the coverage of this previous work in three essential dimensions. First, all three previous surveys specifically focus on form and content of explanations; however, time granularity of explanations has not been investigated. As AVs are real-time decision-making systems, it is crucial to know how explanations must be delivered from the timing perspective. Furthermore, attention-based transformers, large language models, and vision-language models are now at the forefront of AI applied to AVs technologies, and such approaches have not been explored in the aforementioned surveys. Finally, a classification of XAI approaches applied to AVs from an algorithmic/methodological perspective is also a noteworthy nuance missing in these studies. Consequently, our paper extends the above-mentioned reviews by (1) analyzing the temporal sensitivity of explanations, (2) providing a perspective on emerging XAI paradigms, and (3) presenting an algorithmic/methodological taxonomy of XAI methods for AVs. We present the coverage and comparison of these reviews and our work in Table 1.
5.2 Structure of our survey
Our survey presents a comprehensive overview of XAI methods for AVs. In particular, we present a classification of learning approaches in terms of vision, RL and IL, decision trees, logic, user study, and the most recent paradigm shift - large language and vision-language-based explanations for AVs. By adopting insights from the recent industrial trends, emerging AI technologies, and general regulatory-compliance principles, we further present a conceptual framework for explainable end-to-end autonomous driving and describe essential elements of such a framework. Finally, we present a set of promising XAI approaches by describing the missing pieces in the state of the art. We also propose potential solutions with the goal of bridging the gaps and achieving transparency, trustworthiness, and social acceptance in the next-generation AVs.
5.3 Vision-based explanations
As deep neural networks, often in augmented forms of CNNs, power the vision ability of intelligent vehicles, understanding how CNNs capture real-time image segments that lead to the particular behavior of a vehicle is a key concept to achieving visual explanations. In this regard, explainable CNN architectures have resulted in adjustments to generate visual explanations. Zeiler et al. [106] use deconvolution layers to understand the internal representation of CNNs in their seminal work. Hendricks et al. [107] propose a model concentrating on distinguished properties of objects that explain the rationale for the predicted label. Zhou et al.’s [108] saliency map architecture, class activation map (CAM), highlights the discriminative part of an image to predict the label of the image. Moreover, Selvaraju et al. [109] propose an augmented
Figure 8: End-to-end learning of steering angle commands from an input image. Source: [110].
version of CAM, called Grad-CAM, that highlights the derivative of CNN’s prediction with respect to its input. Further examples of backpropagation-based methods include guided-backpropagation, [111], layer-wise relevance propagation [112, 113], and DeepLift [114]. Babiker and Goebel [115, 116] have also shown that heuristics-based Deep Visual Explanations (DVE) provide a justification for predictions of a CNN.
Similarly, Hofmarcher et al. [118] propose a semantic segmentation model implemented as a pixel-wise classification that explains underlying real-time perception of the environment. They evaluate the performance of their framework on Cityscapes [119], a benchmark dataset for understanding street scenes. The framework outperforms other popular segmentation models such as ENet and SegNet with 59.8 per-class mean intersection over union (IoU) and 84.3 per-category mean IoU. Interpretability of the model is a plus for unexpected behaviors, allowing to debug the driving system and understand the rationales for temporal decisions of a self-driving vehicle.
Kim and Canny [110] use a causal attention model on top of the saliency filtering that indicates which input regions actually affect the steering control. Their experiments are conducted on the driving datasets - Comma.ai [120], Udacity [121], and Hyundai Center of Excellence in Integrated Vehicle Safety Systems and Control (HCE): This model runs for nearly 16 hours to train CNNs end-to-end from images to steering angles and apply causality filtering to find out which parts of images have high influence in predictions (Figure 8). With this approach, the learned framework provides an interpretable visualization of a vehicle’s actions. As an enhancement of this model, Kim et al. [122] provide textual explanations in their further study. They produce “intelligible explanations” on the decisive actions of a self-driving vehicle using an attention-based video-to-text mechanism and introduce a novel dataset, called Berkeley Deep Drive-X (eXplanation) (BDD-X), that contains annotations for textual explanations and descriptions.
Zeng et al.’s [123] architecture learns to drive an autonomous vehicle safely by following traffic rules, including interaction with road users, yielding, and traffic signals. They use raw LIDAR data and an HD map to generate interpretable representations as 3D detection of objects, anticipated future trajectories, and cost map visualizations. 3D detection instances provide descriptive information so that the model understands the operational environment. Motion forecasting, measured as L1 and L2 distances, explains
Table 2: Explainable vision-based studies for AVs
@@ 5.3 Vision-based explanations
Echterhoff et al., [146], 2024 Leveraging concept bottlenecks as visual features for predicting control command and explanations of vehicle and human behavior GPT 3.5, Longformer Visual and textual All groups

### v5

@@ 3.1 The need for explanations in AVs
The need for explanations in autonomous driving arises from fundamental issues, established regulations and standards covered in previous subsections, and cross-disciplinary views and opinions of society. At the highest level, the necessity of explanations for AVs can be summarized in terms of four perspectives:
@@
Overall, we can conclude that the explainability of autonomous driving systems is an expectation and a requirement from a multidisciplinary point of view.
3.2 Potential benefits of explanations for AVs
@@
  • Human-Centered Design: Getting the end users’ inputs, opinions, and anticipations on the design and development of the semi or fully AVs can help with the acceptance of this technology by the general community [51].
  • Trustworthiness: Algorithmic assurance can build trust in human-autonomous system relationships [52].
  • Traceability: Explainable intelligent driving systems can help forensic analysts and system auditors understand the entire decision-making process of an autonomous car during the journey via a post-trip analysis.
  • Transparency and accountability: Explanations can help achieve accountability, which can resolve the potential liability and responsibility gaps in foreseeable post-accident investigations with the involvement of autonomous cars as described by Burton et al. [29]. For example, Mercedes-Benz has recently taken a promising step forward and announced that the corporation will take legal responsibility for any accidents that their self-driving systems are engaged in [53]. Mercedes’s declaration of legal culpability is a significant milestone toward the accountability of AVs technology.
@@ 3.3 Explanation recipients in AVs
The details, types, and delivery of explanations vary in accordance with users’ identities, technical background knowledge in autonomous driving, and their various functional and cognitive abilities [45, 54]. For instance, a user having little technical expertise on how AVs operate may be satisfied with a simple explanation of a relevant decision/outcome. However, an autonomous systems engineer will need more informative explanations to understand the current functionalities of the car, with the motivation to appropriately “debug” the existing driving system as required. Therefore, the use of domain knowledge and expertise of the explainee is essential to provide pertinent, sufficiently informative, and intelligible explanations [55, 56]. Motivated by a target audience definition of [45] and [57], we can distinguish four groups of the stakeholders in autonomous driving, namely Group 1 - Road users, Group 2 - AVs developers, Group 3 - Regulators and insurers, and Group 4 - Executive management of automobile companies. Figure 3 provides the identity of such stakeholders and their positions in the relevant classification.
@@ 3.4 How to deliver explanations in AVs?
As explainees are classified based on their domain knowledge and needs, explanations and their design and evaluation techniques also vary depending on the context and knowledge of the category of explainees. In fact, explanation construction is one of the major challenges in current XAI research. Zablocki et al. [58] define four “W” questions in XAI-based autonomous driving: 1) Who needs explanations? 2) Why are explanations needed? 3) What kind of explanations can be generated? and 4) When should explanations be delivered? In general, explanations in AI can be distinguished
@@
based on their derivation category and classification. Some of the early practical studies have applied explanations to automated collaborative filtering systems [59] and knowledge-intensive case-based reasoning [60]. Another empirical approach attempted to derive explanations based on some intelligibility types [61] and used “why,” “why not,” “what if,” and “how to” type explanations for causality filtering. Furthermore, Liao et al. [62] have interviewed twenty user-interface and design practitioners working in different areas of AI to understand users’ explanatory requirements. By doing so, they have attempted to find the gaps in the interviewers’ products and developed a question bank: the authors represent users’ needs as questions so that users may potentially ask about the outcomes produced by an AI system. Overall, the stakeholder needs-based explanation design can be viewed as one of the promising approaches.
Another popular approach to producing explanations is based on using psychological tools from formal theories, according to the literature review of [63]. Depending on the context and addressee, both explanation derivation methods confirm their usefulness. These explanation generation approaches can find alignment in their application in autonomous driving; since autonomous driving involves people with diverse backgrounds in society, relevant XAI design needs inherent adjustments to the context problem.
@@
Except for informational content, the effective communication of explanations is also a key factor for good human-machine teaming. In general, the conveyance of explanations to end users is realized through a user interface (UX) or a human-machine interface (HMI) [64]. For instance, an HMI may be an interface to alert the human driver to take over the control of an vehicle in an emergent situation. Other potential examples are display monitors, sound, light signals, and vibrotactile technology that explain the vehicle’s decision-making intentions and bring situation awareness for people in the loop, as shown in Schneider et al.’s work [65].
4 XAI for autonomous driving: A survey
@@ 4.1 Previous surveys on XAI for AVs
There exist reviews on XAI for AVs, which provide valuable insights from various perspectives. These studies expose a variety of approaches, from a universal look to an algorithmic point of view. In this sense, the first noteworthy review on XAI for AVs is
TABLE I: Studies on visual explanations for AVs
@@
Echterhoff et al., [91], 2024 Leveraging concept bottlenecks as visual features for predicting control command and explanations of vehicle and human behavior Longformer, GPT 3.5 Visual and textual All groups
@@
Omeiza et al.’s work [45]. They study the need for/role of explanations for autonomous driving, and focus on legal requirements, standards, and consumer expectations for the design and development of explainable autonomous driving systems. This provides their basis to present a conceptual XAI framework for modular autonomous driving. In further work, Zablocki et al. [58] present a detailed overview of end-to-end vision-based autonomous driving systems and describe explainability hurdles for AVs from an ML perspective. Finally, in very recent work, Kuznietsov et al. [93] present a systematic review of XAI techniques for modular and end-to-end autonomous driving by focusing on how such techniques improve safety and user trust. In this regard, they propose the SafeX framework for modular autonomous driving by integrating user interface, safety, and explainability.
Our study, as complement, extends the coverage of this previous work in three essential dimensions. First, all three previous surveys specifically focus on form and content of explanations; however, time granularity of explanations has not been investigated. As AVs are real-time decision-making systems, it is crucial to know how explanations must be delivered from the timing perspective. Furthermore, attention-based transformers, large language models, and vision-language models are now at the forefront of AI applied to AVs technologies, and such approaches have not been explored in the aforementioned surveys. Finally, a classification of XAI approaches applied to AVs from an algorithmic/methodological perspective is also a noteworthy nuance missing in these studies. Consequently, our paper extends the above-mentioned reviews by (1) analyzing the temporal sensitivity of explanations, (2) providing a perspective on emerging XAI paradigms, and (3) presenting an methodological taxonomy of XAI methods for AVs.
4.2 Structure of our survey
Our survey presents a comprehensive overview of XAI methods for AVs. In particular, we present a classification of approaches in terms of vision, reinforcement learning, imitation learning, decision trees, logic, user study, and the most recent paradigm
shift - large language and vision-language-based explanations for AVs. By adopting insights from the recent industrial trends, emerging AI technologies, and general regulatory-compliance principles, we further present a conceptual framework for explainable end-to-end autonomous driving and describe its essential elements. Finally, we present a set of promising XAI approaches by describing the missing pieces in the state of the art and present potential solutions with the goal of bridging the gaps and achieving transparency and social acceptance in the next-generation AVs.
4.3 Visual explanations
As deep neural networks, often in augmented forms of CNNs, power the vision ability of intelligent vehicles, understanding how CNNs capture real-time image segments that lead to the particular behavior of a vehicle is a key concept to achieving visual explanations. In this regard, explainable CNN architectures have resulted in adjustments to generate visual explanations. Zeiler et al. [94] use deconvolution layers to understand the internal representation of CNNs in their seminal work. Hendricks et al. [95] propose a model concentrating on distinguished properties of objects that explain the rationale for the predicted label. Zhou et al.’s [96] saliency map architecture, class activation map (CAM), highlights the discriminative part of an image to predict the label of the image. Moreover, Selvaraju et al. [97] propose an augmented version of CAM, called Grad-CAM, that highlights the derivative of CNN’s prediction with respect to its input. Further examples of backpropagation-based methods include guided-backpropagation, [98], layer-wise relevance propagation [99, 100], and DeepLift [101]. Babiker and Goebel [102, 103] have also shown that heuristics-based Deep Visual Explanations (DVE) provide a justification for predictions of a CNN.
Hofmarcher et al. [69] propose a semantic segmentation model implemented as a pixel-wise classification that explains underlying real-time perception of the environment. They evaluate the performance of their framework on Cityscapes [104], a benchmark dataset for understanding street scenes. The framework outperforms other popular segmentation models such as ENet and SegNet with 59.8 per-class mean intersection over union (IoU) and 84.3 per-category mean IoU. Interpretability of the model is a plus for unexpected behaviors, allowing to debug the driving system and understand the rationales for temporal decisions of a self-driving vehicle.
Kim and Canny [67] use a causal attention model on top of the saliency filtering that indicates which input regions actually affect the steering control. Their experiments are conducted on the driving datasets - Comma.ai [105], Udacity [106], and Hyundai Center of Excellence in Integrated Vehicle Safety Systems and Control (HCE): This model runs for nearly 16 hours to train CNNs end-to-end from images to steering angles and apply causality filtering to find out which parts of images have high influence in predictions. With this approach, the learned framework provides an interpretable visualization of a vehicle’s actions. As an enhancement of this model, Kim et al. [68] provide textual explanations in their further study. They produce “intelligible explanations” on the decisive actions of a self-driving vehicle using an attention-based video-to-text mechanism and introduce a novel dataset, called Berkeley Deep Drive-X (eXplanation) (BDD-X), that contains annotations for textual explanations and descriptions.